Subject: APO 3.5m schedule science operations resume
From: Ed Turner
Submitted: Sun, 1 Sep 2002 13:29:14 -0400 (EDT)
Message number: 602
(previous: 601,
next: 605
up: Index)
Following successful completion of on-sky recommissioning of the telescope
and instruments last night, scheduled science operations will resume tonight,
Sunday Sept 1. This concludes an extremely productive shutdown that
required exceptionally hard work and long hours for over 6 weeks by the
APO staff; we should all be very grateful for their efforts and
dedication.
SPICam and the NA2 guider are both inoperative due to problems with their
mechanical cooling systems. They will be returned to service as soon as
possible, depending primarily on the response time of the relevant
vendors. The Echelle is currently warm but is expected to be ready for
use again by Wednesday night.
Ed Turner
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 602 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO